# UNB StepUP: A footStep database for gait analysis and recognition using Underfoot Pressure

### v2

@@ Methods
Following data collection, each participant’s experiment included twelve 30-second balance recordings (three recordings per footwear type) and sixteen 90-second walking recordings (four recordings for each footwear condition), stored as 3D tensors with dimensions of nominally 3000 frames $\times$ 720 pixels $\times$ 240 pixels and 9000 frames $\times$ 720 pixels $\times$ 240 pixels, respectively. It should be noted that the duration of trials may differ slightly by a few frames (typically less than 100 frames, equivalent to about one second) as a result of minor delays when beginning or ending the recordings. The recordings consist of unprocessed, trial-by-trial data capturing continuous standing or walking on the pressure-sensitive tiles. The raw pressure recordings were processed to offer a user-friendly labeled data format suitable for both gait analysis and gait recognition, ensuring quality and consistency. The subsequent sections provide a more detailed explanation of the preprocessing steps for the balance and walking trials.
@@
For the walking trials, individual steps were identified both spatially and temporally within the raw pressure data to enable analysis on a step-by-step basis. Object detection involves determining an object’s location and outline, capturing both its center and boundary. It also includes classifying the object, which, in this context, determines whether or not it is a footstep. Footstep detection and tracking in this study were achieved through simple yet efficient techniques, which present possibilities for real-time online implementation during system operation. Starting from the first frame, the regions of activity on the tiles at each time point of the recording, potentially indicating contact from one or both feet, were identified using a simple connected pixels object detection technique. Specifically, the $720\times 240$ frames of the recording were first converted to binary images using a threshold of 10 kPa. Following this, the frames underwent dilation using a circular structuring element with a radius of 4 and erosion with a circular element of radius 2. Subsequently, pixels that were 2-connected (i.e., separated by no more than two orthogonal hops) were clustered to form objects. In this research, objects whose centroids were within a specific distance limit, generally set at 20 pixels, were combined since they frequently represented the heel and forefoot of a shoe or foot with a high arch. Adjustments to this distance were made to accommodate certain special sole types, including stiletto high heels.
After identifying the bounding boxes for objects within a frame, SORT (Simple Online and Realtime Tracking) [16] was used to track these objects across subsequent frames. SORT employs Kalman filtering to predict object locations by integrating a linear motion model with prior positions. It connects bounding boxes belonging to the same object over time by assessing the overlap between these predicted states and the actual observed positions. This method of footstep extraction produced 3D bounding boxes characterized by dimensions (time, height, width), or ($t,y,x$), for each step, where $x$ and $y$ depend on the size and orientation of the participant’s foot, and $t$ changes according to walking speed. In addition, these bounding boxes are supplied as footstep metadata. Since each pass over the tiles records several consecutive steps (typically 4-6 steps), they enable the computation of various spatiotemporal gait parameters. This encompasses important metrics such as gait speed, step length, step width, stance time, swing time, step angle, and walking path, among others.
@@
To enable more advanced analytic methods, including machine learning and deep learning, it is recommended to store footstep data in tensors, which are essentially multidimensional arrays. Since the shape of each footstep varies due to factors such as foot size, foot angle, and walking speed, padding techniques can be employed to standardize spatial and/or temporal dimensions across different steps. Applying zero padding helps preserve information related to direct factors, such as foot size and walking speed, within the dataset, which may provide significant information for classification [17]. Although these methods enable the arrangement of 3D footstep tensors into an array (i.e., higher-dimensional tensors), they do not facilitate the comparison of plantar pressure in the same foot areas, during gait events, or even at consistent pressure scale levels. In the literature on gait analysis, normalization techniques have been applied in three distinct domains: spatial normalization (e.g., rotating for the foot progression angle and scaling for foot size), temporal normalization (e.g., interpolating to either 100 or 101 frames to represent the full range from $0\%$ to $100\%$ of the gait cycle), and amplitude normalization (e.g., scaling according to body mass or total foot pressure). Although these methods allow for comparisons of footsteps in both spatial and temporal domains, crucial information about subject characteristics might be lost. Consequently, it has been shown that merging various preprocessing pipelines may enhance the efficacy of footstep recognition systems [18].
Given the absence of a universally accepted preprocessing pipeline suitable for all deep learning models and their specific classification or recognition tasks [17], the UNB StepUP-P150 dataset provides two different versions of the extracted footsteps as examples, each demonstrating a separate preprocessing approach. Together with this, a custom Python script is made available, allowing researchers to generate various other configurations of preprocessing pipelines suited to their unique research needs (refer to Table 3 for the list of available method options).
@@
  • Pipeline 2: The second pipeline builds upon the first by incorporating two additional components: (5) spatial resizing and (6) amplitude normalization using the mean GRF. To normalize differences in foot size across participants, the sole dimensions were resized to a common size of 70 pixels in length and 25 pixels in width, corresponding to a size of $35\times 12.5$ cm for the original sensor resolution. The original dimensions of the sole, both length and width, were determined from the aligned footprints by counting the maximum number of active sensors (those with pressure exceeding 10 kPa) present along the two aligned axes during the footstep. To achieve the required dimensions, the pressure maps were resized via nearest neighbour interpolation, and samples were spatially zero-padded to $75\times 40$ pixels for consistency with the other version of the footsteps. To normalize amplitude, the sensor data for each step was subsequently rescaled by dividing by the step’s average GRF. Compared to the initial pipeline, this approach further separates information related to foot size and body weight from the footstep measurements. The initial measurements for sole width, sole length, and average GRF are accessible in the Length, Width, and MeanGRF metadata fields.
@@
The process of data labeling (i.e. data annotation) involves assigning target characteristics to training data, enabling statistical analysis or use in machine learning model development. In this dataset, labels were derived algorithmically from collected footsteps and subsequently confirmed through manual inspection.
@@
  • Side: The left and right labels were determined through a pixel counting method outlined by MacDonald et al. [19]. This method involves counting the number of activated sensors beneath various parts of the foot for use in classification.
@@
  • Orientation: The orientation of each footstep, corresponding to the participant’s direction of walking during each pass across the tiles, was determined by analyzing the COP trajectory of the foot. Foot COP is determined by calculating the average position of the foot in each frame, weighted according to the pressure intensities. To determine the walking direction, the position of the anteroposterior (AP) COP in the initial portion of the footstep was compared with its position in the latter part. A value of 1 indicates walking from Tiles 1 and 2 towards Tiles 11 and 12, while 0 indicates the other direction.
@@
  • Incomplete: Incomplete footsteps were defined as footsteps that fell partially outside of the tile grid or outside of the 90-second recording. These were detected by flagging footsteps with a starting time of $t=0$ or an end time of $t=90$, as well as footsteps that fell near the boundaries of the sensor platform and had a small area. Specifically, footsteps with several active sensors that yielded values smaller than three scaled median absolute deviations (MAD) below the median footstep in the trial were flagged.
@@
  • Standing: Standing footsteps, observed during the slow-to-stop trials, were also identified by analysis of the COP trajectory of the foot. Linear regression using a least-squares method estimated the slope of the AP COP for each footstep. Footsteps with a slope exceeding three MAD from the median footstep in the trial were identified as standing footsteps.
@@
As outlined in the prior section on Data Processing, metadata labels were identified through automated algorithms. Although these algorithms reached impressive accuracy levels (such as $99.7\%$ for identifying left and right footsteps [19], concerning the Side metadata field), the necessity for perfectly accurate labels in benchmark datasets is crucial. Therefore, a minimum of two research team members performed a manual/visual inspection of each footstep and its associated metadata within the dataset. A multi-stage manual inspection of the extracted footsteps was conducted, where any errors identified in the initial review were addressed before starting the second review by a different evaluator. The evaluators conducted assessments of each footstep by visually inspecting the raw data for each pass across the tiles (Fig. 9(a)), along with processed versions of the data for each segmented step (Fig. 9(b)). This included analyzing gait features such as aligned foot peak pressure images, vertical ground reaction force (GRF) profiles, foot center of pressure (COP) trajectories (Fig. 9(b)), as well as reviewing associated video recordings (captured by Camera 7, positioned at a 90° viewing angle, perpendicular to the main axis of the tile grid; Fig. 9(c)). These inspection phases were also used to identify and eliminate any overlooked data corruption, hardware malfunctions, or protocol complications during data collection.
@@
Specifically, the assessors recorded the errors in spreadsheets designated for each trial and merged the necessary corrections into the dataset where applicable. Noted errors fell into two primary categories: those concerning bounding boxes and those related to metadata. An interactive tool was developed to enable evaluators to manually adjust the size of spatiotemporal bounding boxes with inaccurate dimensions. Common bounding box errors included (1) bounding boxes that were too small spatiotemporally to contain a complete footstep, often occurring with high-arch shoes where the algorithm identified just a portion of the pressure profile, such as the heel or toe area; (2) bounding boxes that were too large, encroaching on the spatiotemporal area of another footstep, commonly observed in the W2 trials when participants slightly shuffled their feet during the slow-to-stop walking maneuver; and (3) missing boxes, especially for footsteps that occurred mostly outside the tile grid (e.g., a footstep that fell only 20% within the grid, deemed an incomplete footstep). When common errors were noted in the auto-labeling, corrections were implemented programmatically whenever feasible. Errors most commonly occurred (1) in the Side metadata field, where a left label was incorrectly matched with a right footstep or vice versa, especially for invalid footsteps; (2) in the Standing metadata field, where certain walking behaviours, such as shuffling during a slow-to-stop maneuver, resulted in inconsistent label assignment by the algorithm; and (3) in the Incomplete metadata field, where partial footsteps were not flagged as incomplete steps.

### v3

@@ Methods
A total of 28 raw pressure data recordings were captured for each participant, twelve 30-second balance recordings (three recordings for each footwear condition) and sixteen 90-second walking recordings (four recordings for each footwear condition). Each raw data recording is stored as a 3D tensor with dimensions of nominally 3000 frames $\times$ 720 pixels $\times$ 240 pixels and 9000 frames $\times$ 720 pixels $\times$ 240 pixels, for the 30-second and 90-second trials respectively. It should be noted that the duration of trials may differ slightly by a few frames (typically less than 100 frames, equivalent to about one second) as a result of minor delays when beginning or ending the recordings. The recordings consist of unprocessed, trial-by-trial data capturing continuous standing or walking on the pressure-sensitive tiles. These raw pressure recordings were processed to offer a user-friendly labeled data format suitable for both gait analysis and gait recognition, ensuring quality and consistency. The subsequent sections provide a more detailed explanation of the preprocessing steps for the balance and walking trials.
@@
For the walking trials, individual steps were identified both spatially and temporally within the raw pressure data to enable analysis on a step-by-step basis. Footstep detection and tracking in this study were achieved through simple yet efficient techniques, which present possibilities for real-time online implementation during system operation. Starting from the first frame, the regions of activity on the tiles at each time point of the recording, potentially indicating contact from one or both feet, were identified using a simple connected pixels object detection technique. Specifically, the $720\times 240$ frames of the recording were first converted to binary images using a threshold of 10 kPa. The resulting binary frames were then processed with morphological operations: dilation using a circular structuring element with a radius of 4, followed by erosion with a circular element of radius 2. Subsequently, pixels that were 2-connected (i.e., separated by no more than two orthogonal hops) were clustered to form objects. In this research, objects whose centroids were within a specified distance (e.g., 20 pixels) were combined since they frequently represented the heel and forefoot of a shoe or foot with a high arch. Adjustments to this distance were made as needed to accommodate certain special sole types, including stiletto high heels.
After identifying the bounding boxes for objects within a frame, SORT (Simple Online and Realtime Tracking) [16] was used to track these objects across subsequent frames. SORT uses a Kalman filter combined with a linear motion model to predict object locations based on prior positions, linking bounding boxes over time by assessing the overlap between predicted and observed positions. This method of footstep extraction produced 3D bounding boxes with dimensions (time, height, width), or ($t,y,x$), for each step, where $x$ and $y$ depend on the size and orientation of the participant’s foot, and $t$ changes according to walking speed. These bounding boxes are supplied as footstep metadata. Since each pass over the tiles records several consecutive steps, they enable the computation of various spatiotemporal gait parameters including gait speed, step length, step width, stance time, swing time, step angle, and walking path, among others.
@@
To enable more advanced analytic methods, including machine learning and deep learning, it is recommended to store footstep data in tensors, which are essentially multidimensional arrays. Since the shape of each footstep varies due to factors such as foot size, foot angle, and walking speed, padding techniques can be employed to standardize spatial and/or temporal dimensions across different steps. Applying zero padding helps preserve information related to direct factors, such as foot size and walking speed, within the dataset, which may provide significant information for classification [17]. While these methods allow the 3D footstep tensors to be organized into arrays (i.e., higher-dimensional tensors), they do not support meaningful comparisons of plantar pressure across consistent foot regions, gait events, or pressure intensity scales. In the literature on gait analysis, normalization techniques have been applied in three distinct domains: spatial normalization (e.g., rotating for the foot progression angle and scaling for foot size), temporal normalization (e.g., interpolating to either 100 or 101 frames to represent the full range from $0\%$ to $100\%$ of the gait cycle), and amplitude normalization (e.g., scaling according to body mass or total foot pressure). Although these methods allow for comparisons of footsteps in both spatial and temporal domains, crucial information about subject characteristics might be lost. Consequently, it has been shown that merging various preprocessing pipelines may enhance the efficacy of footstep recognition systems [18].
Given the absence of a universally accepted preprocessing pipeline suitable for all deep learning models and their specific classification or recognition tasks [17], the UNB StepUP-P150 dataset provides two different versions of the extracted footsteps as examples, each demonstrating a separate preprocessing approach. Together with this, a custom Python script is made available, allowing researchers to generate various other configurations of preprocessing pipelines suited to their unique research needs (refer to Table 3 for the list of provided options).
@@
  • Pipeline 2: The second pipeline builds upon the first by incorporating two additional components: (5) spatial resizing and (6) amplitude normalization using the mean total pressure. To normalize differences in foot size across participants, the sole dimensions were resized to a common size of 70 pixels in length and 25 pixels in width, corresponding to a size of $35\times 12.5$ cm for the original sensor resolution. The original dimensions of the sole, representing the length and width of the foot’s contact area with the sensor grid, were determined from the aligned footprints by counting the maximum number of active sensors (those with pressure exceeding 10 kPa) present along the two aligned axes during the footstep. To achieve the required dimensions, the pressure maps were resized via nearest neighbour interpolation, and samples were spatially zero-padded to $75\times 40$ pixels for consistency with the other version of the footsteps. To normalize amplitude, the sensor data for each step was subsequently rescaled by dividing by the time-averaged sum of all sensor readings (i.e., the average total pressure). Compared to the initial pipeline, this approach further separates information related to foot size and body weight from the footstep measurements. The initial measurements for sole width, sole length, and average pressure are accessible in the Length, Width, and MeanPressure metadata fields.
@@
The process of data labeling (i.e. data annotation) involves assigning target characteristics to training data, enabling statistical analysis or use in machine learning model development. In this dataset, labels were derived algorithmically from collected footsteps and subsequently confirmed through manual inspection (see details in the Manual Data Inspection section).
@@
  • Side: The left and right foot labels were determined through a pixel counting method outlined by MacDonald et al. [19]. This method involves counting the number of activated sensors beneath different regions of the foot.
@@
  • Orientation: The orientation of each footstep, corresponding to the participant’s direction of walking during each pass across the tiles, was determined by analyzing the COP trajectory of the foot (see Equation (2) in the Technical Validation section). To determine the walking direction, the position of the anteroposterior (AP) COP in the beginning of the footstep was compared with its position at the end. A value of 1 indicates walking from Tiles 1 and 2 towards Tiles 11 and 12, while 0 indicates the opposite.
@@
  • Incomplete: Incomplete footsteps were defined as footsteps that fell partially outside of the tile grid or outside of the 90-second recording. These were detected by flagging footsteps with a starting time of $t=0$ or an end time of $t=90$, as well as footsteps that fell near the boundaries of the sensor platform and had a small area. Specifically, footsteps with a total number of activated sensors smaller than three scaled median absolute deviations (MAD) below the median footstep in the trial were flagged.
@@
  • Standing: Standing footsteps, observed during the slow-to-stop trials, were also identified by analysis of the COP trajectory of the foot. Linear least-squares regression was used to estimate the slope of the AP COP for each footstep, and footsteps with a slope exceeding three MAD from the median footstep in the trial were identified as standing footsteps.
Data Records
The StepUP-P150 dataset is accessible for download on the Federated Research Data Repository (FRDR) [20] (https://doi.org/10.20383/103.01285) and contains the raw recordings, two variants of extracted and preprocessed footstep data, and detailed metadata for each individual footstep. The files are organized as shown in Fig. 9. Located at the top level, the spreadsheet named ‘participant_metadata.csv’ contains demographic information, anthropometric measurements, types of footwear, and other possible influences on the gait patterns of each for the 150 participants. More comprehensive information regarding the metadata fields is available in Table 4. For convenience, both Python-compatible and MATLAB-compatible versions of the footstep data are provided, organized into the top-level folders ‘py’ and ‘mat’, respectively. For each version, the pressure recordings are structured into 150 folders, with each corresponding to a single participant. Each folder is named using the pattern XXX, with ‘XXX’ representing a unique, randomly assigned ID that encompasses all participants (001-150). Within each participant’s folder, four (third-level) sub-folders represent the various footwear conditions: ‘BF’, ‘ST’, ‘P1’, and ‘P2’. Inside these, there are seven further (fourth-level) sub-subfolders designated for distinct standing balance and walking experiment trials: ‘S1’, ‘S2’, ‘S3’, ‘W1’, ‘W2’, ‘W3’, or ‘W4’ (refer to Fig. 4 and the Experimental Protocol section for more information).
For each of the three standing balance trials (S1, S2, and S3), there are two corresponding files. The ‘trial.{npz,mat}’ file houses the raw, unprocessed 30-second recording from the trial, represented as a 3D tensor with dimensions of nominally $3000\times 720\times 240$ (frames, height, width; note that the number of frames may vary slightly). The processed version of the recording is available as ‘preprocessed.{npz,mat}’, representing a 3D tensor with dimensions of $3000\times 180\times 180$ (frames, height, width). As noted, these files are available in both NumPy (.npz) and MATLAB (.mat) file formats, each of which uses a dictionary-like structure for variable storage, with the tensors accessible under the top-level key arr_0.
Unlike the balance trials, each folder for the walking trials (W1, W2, W3, and W4) includes four files. The raw, unprocessed 90-second pressure data is available in ‘trial.{npz,mat}’, containing a 3D tensor with dimensions of $9000\times 720\times 240$ (frames, height, width; again, the number of frames may vary). The footsteps were extracted using two distinct preprocessing methods and are provided as ‘pipeline_1.{npz,mat}’ and ‘pipeline_2.{npz,mat}’, which are explained in the Footstep Normalization section. Each tensor is 4D with dimensions $n_{footsteps}\times 101\times 75\times 40$ (samples, frames, height, width), where $n_{footsteps}$ specifies the number of footsteps in a given trial. These tensors are all accessible via the key arr_0. Each individual footstep’s metadata is provided in a file named ‘metadata.csv’, containing details such as the footstep’s spatiotemporal location in the initial recording, classification labels (e.g., left or right foot, walking direction, outlier), and various parameters obtained during preprocessing (e.g., rotation angle, length and width of the foot’s contact area, mean pressure). Table 5 lists the metadata fields and their explanations.
Technical Validation
@@
As outlined in the prior section on Data Processing, metadata labels were identified through automated algorithms. Although these algorithms reached impressive accuracy levels (such as $99.7\%$ for identifying left and right footsteps [19] for the Side metadata field), the necessity for perfectly accurate labels in benchmark datasets is crucial. Therefore, a minimum of two research team members performed a manual/visual inspection of each footstep and its associated metadata within the dataset. A multi-stage manual inspection of the extracted footsteps was conducted, where any errors identified in the initial review were addressed before starting the second review by a different evaluator. The evaluators conducted assessments of each footstep by visually inspecting the raw data for each pass across the tiles (Fig. 10(a)), along with processed versions of the data for each segmented step (Fig. 10(b)). This included analyzing gait features such as aligned foot peak pressure images, vertical ground reaction force (GRF) profiles, foot center of pressure (COP) trajectories (Fig. 10(b)), as well as reviewing associated video recordings (captured by Camera 7, positioned at a 90° viewing angle, perpendicular to the main axis of the tile grid; Fig. 10(c)). These inspection phases were also used to identify and eliminate any overlooked data corruption, hardware malfunctions, or protocol complications during data collection.
@@
Specifically, the assessors recorded the errors in spreadsheets designated for each trial and merged the necessary corrections into the dataset where applicable. Errors fell into two primary categories: those concerning bounding boxes and those related to metadata. An interactive tool was developed to enable evaluators to manually adjust the size of spatiotemporal bounding boxes with inaccurate dimensions. Common bounding box errors included (1) bounding boxes that were too small spatiotemporally to contain a complete footstep, often occurring with high-arch shoes where the algorithm identified just a portion of the pressure profile, such as the heel or toe area; (2) bounding boxes that were too large, encroaching on the spatiotemporal area of another footstep, commonly observed in the W2 trials when participants slightly shuffled their feet during the slow-to-stop walking maneuver; and (3) missing boxes, especially for footsteps that occurred mostly outside the tile grid (e.g., a footstep that fell only 20% within the grid, deemed an incomplete footstep). When common errors were noted in the auto-labeling, corrections were implemented programmatically whenever feasible. Errors most commonly occurred (1) in the Side metadata field, where a left label was incorrectly matched with a right footstep or vice versa, especially for incomplete footsteps; (2) in the Standing metadata field, where certain walking behaviours, such as shuffling during a slow-to-stop maneuver, resulted in inconsistent label assignment by the algorithm; and (3) in the Incomplete metadata field, where partial footsteps were not flagged as incomplete steps.